OpenStudy (anonymous):

( (5/-5)k ) + (-3/7) = -7 - ( (-7/4)k )

OpenStudy (anonymous):

File is attached.. couldn't get this one the first time too! I need to know how to solve this problem. :/

OpenStudy (anonymous):

k=72/77

OpenStudy (anonymous):

Incorrect.

OpenStudy (anonymous):

184/77 ?

OpenStudy (anonymous):

incorrect

OpenStudy (anonymous):

k=184/77

OpenStudy (anonymous):

it says incorrect.. :/

OpenStudy (anonymous):

13/2=6.5

OpenStudy (anonymous):

I missed the screenshot .

OpenStudy (anonymous):

thats incorrect too?

OpenStudy (anonymous):

Alright so let's start by simplifying these signs. ( (5/-5)k ) + (-3/7) = 7 - ( (-1/7)k ) -k - 3/7 = 7 + 1/7k -k - 3/7k = 7 + 3/7 -8/7k = 52/7 <--- Changed 7 into 49 /7 and added 3/7 k = -52/8 <---- Simplify top at bottom by 4 k = -13/ 2

OpenStudy (anonymous):

Sorry third line was -k - 1/7k = 7 + 3/7

OpenStudy (anonymous):

ohh ok! i see now thank you!!!

OpenStudy (anonymous):

No problem, glad I could help :D